# Dual Prototyping with Domain and Class Prototypes for Affective Brain-Computer Interface in Unseen Target Conditions

### v1

@@ Abstract
EEG signals have emerged as a powerful tool in affective brain–computer interfaces, playing a crucial role in emotion recognition. However, current deep transfer learning-based methods for EEG recognition face challenges due to the reliance of both source and target data in model learning, which significantly affect model performance and generalization. To overcome this limitation, we propose a novel framework (PL-DCP) and introduce the concepts of feature disentanglement and prototype inference. The dual prototyping mechanism incorporates both domain and class prototypes: domain prototypes capture individual variations across subjects, while class prototypes represent the ideal class distributions within their respective domains. Importantly, the proposed PL-DCP framework operates exclusively with source data during training, meaning that target data remains completely unseen throughout the entire process. To address label noise, we employ a pairwise learning strategy that encodes proximity relationships between sample pairs, effectively reducing the influence of mislabeled data. Experimental validation on the SEED and SEED-IV datasets demonstrates that PL-DCP, despite not utilizing target data during training, achieves performance comparable to deep transfer learning methods that require both source and target data. This highlights the potential of PL-DCP as an effective and robust approach for EEG-based emotion recognition.
Index Terms:
Electroencephalography; Prototype Inference; Transfer learning; Unseen target; Emotion Recognition.I Introduction
Emotion, as a fundamental physiological signal, plays a critical role in human communication. Extensive research has established a close link between emotional states and mental health disorders [1]. Accurately recognizing emotions is crucial for daily life, psychological health management, and human-computer interaction. Recently, electroencephalography (EEG)-based emotion recognition has become a focal point for understanding and modulating human emotions, attracting growing interest from researchers in affective computing [2, 3]. This trend is further fueled by advancements in machine learning and deep learning, which have enabled the development of increasingly efficient and accurate models for emotion recognition [4]. Despite this progress, two critical challenges remain that limit the practical application of EEG-based emotion recognition systems. First, EEG signals are highly individualized, and current models rely heavily on target domain data (transfer learning models), which reduces their adaptability in real-world settings. How can we create emotion recognition models that accommodate substantial individual variability without requiring target-specific data, thereby improving real-world usability? Second, label noise continues to be a pressing issue, affecting model reliability. Developing models with increased resilience to label noise is essential for enhancing the robustness of emotion recognition systems. Addressing these challenges is crucial for advancing the field and enabling more adaptable, accurate, and reliable emotion recognition technologies.
@@
Previous studies have emphasized that EEG signals are highly subject-dependent [5, 6], and that the way individuals perceive and express emotions can vary significantly [7]. These individual differences extend to neural processes involved in emotional regulation, further complicating the task of emotion recognition [8]. A fundamental assumption in many machine learning and deep learning methods is that training and testing data share the same feature space and follow the same distribution, thus satisfying the independent and identically distributed (IID) condition. However, individual variability in EEG signals often disrupts this assumption, leading to substantial performance degradation or even model failure when traditional emotion recognition models are applied to new subjects. This variability presents a significant challenge for the effectiveness and generalizability of existing models, underscoring the need for approaches that can adapt to diverse individual EEG patterns without compromising performance.
@@
To address the challenges posed by individual variability in EEG signals, a growing number of researchers are employing transfer learning methodologies, which have shown promising results [9, 10, 11, 12, 13, 14, 15, 16]. Transfer learning accommodates variations in domains, tasks, and data distributions between training and testing phases. By considering the feature distributions of both the source domain (with labeled data and known distribution) and the target domain (with unlabeled data and unknown distribution), transfer learning approach leverages knowledge from the source domain to improve predictive accuracy in the target domain [17]. Given the success of transfer learning in addressing individual differences in EEG signals, most current EEG-based emotion recognition models are developed within a transfer learning framework [18, 12, 19, 13]. On the other hand, transfer learning models generally require simultaneous access to both source and target domain data during training, which necessitates retraining the model before it can adapt to new subjects. This requirement significantly increases the practical costs associated with model deployment, particularly when dealing with large datasets or complex models with numerous parameters. Retraining in these cases can be time-consuming and computationally intensive, posing a challenge for efficient deployment and limiting the scalability of these models in real-world, diverse scenarios. Addressing this limitation is crucial for developing more adaptable and cost-effective EEG-based emotion recognition systems. Therefore, addressing the individual differences in EEG signals while avoiding dependence on target domain data is a crucial step for EEG-based emotion recognition models to advance towards practical applications.
To address the aforementioned issues, we propose a Pairwise Learning Framework with Domain and Class Prototypes (PL-DCP) for EEG-based emotion recognition in unseen target conditions. In this framework, we incorporate not only domain prototypes but also class prototypes to effectively capture diverse feature distributions and achieve improved alignment across data distributions. To mitigate the effects of label noise, classification is formulated as a pairwise learning task, evaluating the relationships between samples and the various prototypes. Importantly, to enhance practicality, the proposed model does not rely on target domain data during training, making it well-suited for real-world applications. The main contributions are summarized below. (1) Dual Prototyping with Domain and Class Prototypes. EEG features are represented through a novel dual prototyping manner (Fig. 1): domain prototype inference (representing the domain to which the sample belongs) and class prototype inference (indicating the emotional class of the sample). Individual variability in EEG signals is thus conceptualized as a feature shift resulting from the interaction between these domain and class prototypes. (2) Independence from Target Domain Data for Improved Practicality. To enhance generalizability and facilitate broader deployment of emotion recognition systems, the model is designed to operate without exposure to target domain data during training. (3) Robust Performance without Target Domain Data. Experimental results show that the proposed PL-DCP achieves performance comparable to, or even surpassing, classical deep transfer learning models, despite the absence of target domain data.
II Related Work
@@
II-A Non-deep transfer learning models
@@
II-B Deep transfer learning models
@@
Non-deep transfer models would be limited in complexity and capacity, constraining their ability to fully meet the practical demands of emotion recognition. With advances in deep learning theories and technologies, deep learning-based transfer algorithms have been introduced, offering enhanced model capabilities in terms of performance and generalizability. These algorithms have been widely applied in EEG-based emotion recognition, and most contemporary models now leverage deep transfer learning.
For example, Tzeng et al. [24] proposed the Deep Domain Confusion (DDC) architecture, a CNN model incorporating an adaptation layer and a domain confusion loss based on MMD. This approach learns representations that are both semantically rich and domain-invariant. Zhang et al. [25] introduced a cross-subject emotion recognition method that utilizes CNNs with DDC. This method constructs an Electrode-Frequency Distribution Map (EFDM) from EEG signals, using a CNN to extract emotion-related features while employing DDC to minimize distribution differences between source and target domains. Jin et al. [26] applied the Domain-Adversarial Neural Network (DANN) framework to EEG-based emotion recognition. DANN eliminates distribution discrepancies between source and target subjects, improving cross-domain robustness. Many recent deep transfer learning methods for EEG emotion recognition have been developed based on the DANN structure, leveraging its capacity for effective domain adaptation. For example, Li et al. [18] introduced the Bi-Domain Adversarial Neural Network (BiDANN), which accounts for asymmetrical emotional responses in the left and right hemispheres of the brain, leveraging neuroscientific insights to improve emotion recognition performance. Li et al. [27] developed the Region-Global Spatial-Temporal Neural Network (R2G-STN), which incorporates neuroscientific insights by considering emotional response variations across different brain regions. Ye et al. [16] introduced a semi-supervised Dual-Stream Self-Attentive Adversarial Graph Contrastive learning framework (DS-AGC) to enhance feature representation in scenarios with limited labeled data. This framework includes a graph contrastive learning method to extract effective graph-based feature representations from multiple EEG channels. Additionally, it incorporates a self-attentive fusion module for feature fusion, sample selection, and emotion recognition, focusing on EEG features more relevant to emotions and identifying data samples in the labeled source domain that are closer to the target domain. These deep transfer learning methods has provided valuable insights and strategies for addressing individual differences in EEG signals and achieving significant results in EEG-based emotion recognition. However, a key challenge remains: the reliance on target domain data during training, which increases practical application costs and limits the scalability of these models in real-world scenarios.
II-C Prototype Learning
The core concept of prototype learning is that each class is represented by a prototype (a feature vector that acts as a central, representative feature for that class). Data points belonging to a specific class are clustered around this prototype, enabling classification by evaluating the proximity or similarity of data points to their respective class prototypes. For example, Snell et al. [28] proposed prototypical networks, which learn a metric space where samples from the same category are clustered around their respective class prototypes. In Pinheiro et al. [29]’s work, prototype representations are computed for each category, and target domain images are classified by comparing their feature representations to these prototypes, assigning the label of the most similar prototype. Ji et al. [30] tackled proposed Semantic-guided Attentive Prototypes Network (SAPNet) framework to address the challenges of extreme imbalance and combinatorial explosion in Human-Object Interaction (HOI) tasks. Liu et al. [31] developed a refined prototypical contrastive learning network for few-shot learning (RPCL-FSL), which combines contrastive learning with few-shot learning in an end-to-end network for enhanced performance in low-data scenarios. Yang et al. [32] introduced the Two-Stream Prototypical Learning Network (TSPLN), which simultaneously considers the quality of support images and their relevance to query images, thereby optimizing the learning of class prototypes. These studies demonstrate that prototype learning is particularly effective in few-shot learning and unsupervised tasks, offering a potential solution by enabling models that do not rely on target domain data. In the application of EEG-based emotion recognition, Zhou et al. [9] proposed a prototype representation-based pairwise learning framework (PR-PL), where sample features interact with prototype features through a bilinear transformation. Experimental results demonstrated that these interaction features with prototype representations can significantly enhance model performance. However, PR-PL considers only class prototypes, assuming that source domain data follow a uniform distribution, a limitation that does not align with real-world variability. Additionally, PR-PL requires both source and target data during training to align sample features, which restricts its practical applicability in scenarios where target data may not be accessible.
@@ III Methodology
The source domain is defined as ={,,,…,, where denotes the number of subjects in the source domain. For each individual subject in the source domain, we have ={,, where denotes the -th sample of -th subject, represents the corresponding emotion label, and is the sample size for the -th subject. The target domain is represented as ={,, where denotes the number of EEG samples in the target domain. For clarity, Table I summarizes the commonly used notations.
III-A Feature Disentanglement
@@
Inspired by the work of Peng et al. [33] and Cai et al. [34], we hypothesize that EEG features involves two types of deep features: domain-invariant class features and class-invariant domain features. The domain-invariant class features capture semantic information about the class to which a sample belongs, while the class-invariant domain features convey the domain or subject-specific information of the sample. Original EEG features can be viewed as an integration of these two types of features.
@@
The distributional differences in EEG signals across subjects can be attributed to variations in the domain features, causing a shift in the distribution of class features. Since EEG classification adheres to a common standard, the class features from different subjects should ideally occupy the same feature space and follow a consistent distribution. Traditional methods assume that all test data come from a single, unified domain, effectively focusing only on class features while overlooking the presence of domain features. This assumption limits model generalization across subjects. Feature extraction methods such as DANN can be interpreted as an attempt to remove the domain-specific components from sample features, thereby aligning the class features across subjects and improving the generalization performance of the model. In this paper, we approach EEG feature extraction with consideration for both domain-specific and class-specific components to improve robustness and generalization across diverse subjects.
@@
We start with a shallow feature extractor to obtain shallow features from the EEG samples . Then, we introduce a class feature disentangler and a domain feature disentangler to disentangle the semantic information within these shallow features, resulting in class features and domain features , expressed as:
@@
To improve the effectiveness of the disentanglers in separating the two types of features, we introduce a domain discriminator and a class discriminator . The domain discriminator is designed to determine the domain of the input features, while the class discriminator ascertains the class of the input features. Our goal is for the domain discriminator to accurately identify the domain of the input when given domain features, while the class discriminator should be unable to identify the class based solely on domain features. This inability to classify based on domain features indicates successful disentanglement, where domain features contain only domain-specific information and no class-related information. Similarly, we aim for class features to contain only class-related information, free from domain-specific information. To achieve this, we draw on ideas from DANN [35]. Before the class features are passed into the domain discriminator and the domain features into the class discriminator, they pass through a Gradient Reversal Layer (GRL) to facilitate adversarial training. We use a binary cross-entropy loss function to optimize the discriminators. The output from each discriminator is first passed through a sigmoid layer to obtain probability values, which are then compared to the true labels. This approach converts the multi-class problem into several independent binary classification tasks. The binary cross-entropy loss function is defined as follows:
@@
where represents the GRL. represents the true class labels of -th sample, represents the class feature of -th sample, and represents the domain feature -th sample. This loss function is optimized to help the class discriminator learn to distinguish class labels accurately, ensuring that the class features retain only class-related information, with minimal interference from domain-specific features. Similarly, for the domain discriminator, we have the following binary cross-entropy loss function:
@@
Here, represents the true domain labels of -th sample. This loss function is optimized to ensure that the domain discriminator accurately identifies the domain-specific information, encouraging the domain features to be disentangled from class-related information. By applying the GRL before the domain features enter the class discriminator, adversarial training is facilitated, further promoting the separation of domain and class features. In the implementation, the shallow feature extractor, domain feature disentangler, and class feature disentangler are all designed as multi-layer perceptrons (MLPs). The disentangled class features and domain features are then utilized in the subsequent prototype inference module. This module learns domain prototypes to represent each domain. It also learns class prototypes to capture each class within those domains.
III-B Prototype Inference

### v2

# PL-DCP: A Pairwise Learning framework with Domain and Class Prototypes for EEG emotion recognition under unseen target conditions

@@ Abstract
Electroencephalogram (EEG) signals serve as a powerful tool in affective Brain-Computer Interfaces (aBCIs) and play a crucial role in affective computing. In recent years, the introduction of deep learning techniques has significantly advanced the development of aBCIs. However, the current emotion recognition methods based on deep transfer learning face the challenge of the dual dependence of the model on source domain and target domain, As well as being affected by label noise, which seriously affects the performance and generalization ability of the model. To overcome this limitation, we proposes a Pairwise Learning framework with Domain and Category Prototypes for EEG emotion recognition under unseen target conditions (PL-DCP), and integrating concepts of feature disentanglement and prototype inference. Here, the feature disentanglement module extracts and decouples the emotional EEG features to form domain features and class features, and further calculates the dual prototype representation. The Domain-pprototype captures the individual variations across subjects, while the class-prototype captures the cross-individual commonality of emotion categories. In addition, the pairwise learning strategy effectively reduces the noise effect caused by wrong labels. The PL-DCP framework conducts a systematic experimental evaluation on the published datasets SEED, SEED-IV and SEED-V, and the accuracy are 82.88%, 65.15% and 61.29%, respectively. The results show that compared with other State-of-the-Art(SOTA) Methods, the PL-DCP model still achieves slightly better performance than the deep transfer learning method that requires both source and target data, although the target domain is completely unseen during the training. This work provides an effective and robust potential solution for emotion recognition. The source code is available at https://github.com/WuCB-BCI/PL_DCP.
keywords:
[…]
1 Introduction
Affective, as a basic psychological state, plays a crucial role in daily life. It not only affects people’s feelings, thoughts and behaviors, but also affects people’s physical and mental health.[1] Extensive research has established a strong link between emotional states and mental health disorders.[2] As a physiological signal, Electroencephalogram (EEG) has the characteristics of high time resolution and objectivity, which provides more direct and objective clues for understanding and evaluating emotional states.[3, 4] Affective computing is a rapidly developing interdisciplinary research field. Accurately recognizing emotions and providing personalized treatment plans are key challenges in the field of emotion computing, which is crucial for daily life, mental health management, human-computer interaction and so on.[5]
In recent years, with the further development of deep learning technology, the field of emotion recognition has also been vigorously developed,[6] and has attracted more and more attention from researchers in different fields such as computer science and psychology. It has become a research hotspot in understanding and recognizing human emotions.[7][8] For example:
Feng et al. [9] designed a hybrid model called ST-GCLSTM, which comprises a spatial-graph convolutional network (SGCN) module and an attention-enhanced bi-directional Long Short-Term Memory (LSTM) module, which can be used to extract representative spatial-temporal features from multiple EEG channels.
Yang et al. [10] proposed spectral-spatial attention alignment multi-source domain adaptation (-MSD), which constructs domain attention to represent affective cognition attributes in spatial and spectral domains and utilizes domain consistent loss to align them between domains.
Zhang et al. [11] introduced both cascade and parallel convolutional recurrent neural network models for precisely identifying human intended movements and instructions, effectively learning the compositional spatio-temporal representations of raw EEG streams.
Despite these advances, two key challenges remain that limit the practical application of EEG-based emotion recognition systems.
(1) EEG signals are highly personalized. Existing transfer learning models rely heavily on target domain data, which makes the model vulnerable to data preferences, thus reducing its adaptability in the real world environment and increasing the cost of model training. Therefore, we should create emotion recognition models under the condition of unseen target domain to adapt to a large number of individual differences, so as to improve real-world usability.
(2) Currently, EEG emotion experiments are basically induced by emotional stimulation material, such as video, pictures, audio, etc. Due to individual physiological factors, subjects may not always respond correctly to emotions. Label noise remains a pressing issue that affects the reliability of the model. It is very important to build an emotion recognition model with the ability to resist label noise to improve the robustness of emotion recognition system. Therefore, addressing these two challenges is crucial to advance the field and achieve more adaptable, accurate and reliable emotion recognition techniques.
@@ 1 Introduction
Previous studies have emphasized that EEG signals are highly subject-dependent,[12][13] and that the way individuals perceive and express emotions can vary significantly [14]. These individual differences extend to neural processes involved in emotional regulation, further complicating the task of emotion recognition. [15] A fundamental assumption in many machine learning and deep learning methods is that training and testing data share the same feature space and follow the same distribution, thus satisfying the independent and identically distributed (IID) condition.[16][17] However, individual variability in EEG signals often disrupts this assumption, leading to substantial performance degradation or even model failure when traditional emotion recognition models are applied to new subjects. This variability presents a significant challenge for the effectiveness and generalizability of existing models, underscoring the need for approaches that can adapt to diverse individual EEG patterns without compromising performance.
@@
To address the challenges posed by individual variability in EEG signals, a growing number of researchers are employing transfer learning methodologies, which have shown promising results.[18, 19, 20, 21, 22, 3] Transfer learning accommodates variations in domains, tasks and data distributions between training and testing phases. By considering the feature distributions of both the source domain (with labeled data and known distribution) and the target domain (with unlabeled data and unknown distribution), transfer learning approach leverages knowledge from the source domain to improve predictive accuracy in the target domain.[23] Given the success of transfer learning in addressing individual differences in EEG signals, most current EEG-based emotion recognition models are developed within a transfer learning framework.[24, 25, 26, 27]
In addition, transfer learning models generally require simultaneous access to both source and target domain data during training, which necessitates retraining the model before it can adapt to new subjects. This requirement significantly increases the practical costs associated with model deployment, particularly when dealing with large datasets or complex models with numerous parameters. Retraining in these cases can be time-consuming and computationally intensive, posing a challenge for efficient deployment and limiting the scalability of these models in real-world, diverse scenarios. Addressing this limitation is crucial for developing more adaptable and cost-effective EEG-based emotion recognition systems. Therefore, addressing the individual differences in EEG signals while avoiding dependence on target domain data is a crucial step for EEG-based emotion recognition models to advance towards practical applications.
In order to solve the above problems, We proposes a Pairwise Learning transfer framework based on Domain and Class Prototype (PL-DCP). In this framework, we propose a domain-class dual feature decoupling mechanism for EEG signals, where individual differences are modeled as the superimposed offset of domain features and class features, and fine-grained feature separation is achieved through semantic structure encoding. Domain prototypes (representing the individual differences of subjects, as shown in Fig.1.(a)) and class prototypes (representing the commonality across subjects’ emotional categories, as shown in Fig.1.(b)) are innovatively constructed, and the sample distribution of the target domain is dynamically calibrated by using prototype similarity matching to alleviate the bottleneck of model generalization caused by individual differences of EEG signals. Alternatively, to reduce the impact of label noise, classification is described as a pairwise learning task that evaluates the relationship between samples and different prototypes. It is worth noting that the target domain data is not needed in the training process, which breaks through the dependence of traditional deep transfer learning on target domain data. Overall, The main contributions are summarized below:
EEG features are represented through a novel dual prototype of domain and class prototypes, enabling individual variability in EEG signals to be conceptualized as feature shifts resulting from interactions between these domains and class prototypes.
Proposed the concepts of feature disentanglement and prototype reasoning, obtained domain features and class features from EEG signals. And it does not need to touch the target domain data during the training process.
We rigorously tested model using public databases. Experimental results show that, in the absence of target domain data, PL-DCP still achieves comparable or even better performance than classical deep transfer learning models. In addition, we perform a thorough analysis of our performance on the model and feature visualization to deepen our understanding of the model and results.
2 Related Work
@@
2.1 Non-deep transfer learning models
@@
2.2 Deep transfer learning models
@@
Non-deep transfer models would be limited in complexity and capacity, constraining their ability to fully meet the practical demands of emotion recognition. With advances in deep learning theories and technologies, deep learning-based transfer algorithms have been introduced, offering enhanced model capabilities in terms of performance and generalizability. These algorithms have been widely applied in EEG-based emotion recognition, and most contemporary models now leverage deep transfer learning. For examples,
Lin et al. [32] proposed a novel emotion recognition method based on a novel deep learning model (ERDL). The model fuses graph convolutional neural network (GCNN) and long-short term memories neural networks (LSTM), to extract graph domain features and temporal features.
Ye et al. [3] introduced a semi-supervised Dual-Stream Self-Attentive Adversarial Graph Contrastive learning framework (DS-AGC) to enhance feature representation in scenarios with limited labeled data. The DS-AGC includes a graph contrastive learning method to extract effective graph-based feature representations from multiple EEG channels. Additionally, it incorporates a self-attentive fusion module for feature fusion, sample selection, and emotion recognition.
In recent year, many recent deep transfer learning methods for EEG emotion recognition have been developed based on the Domain-Adversarial Neural Network (DANN[33]) structure, leveraging its capacity for effective domain adaptation. For example,
Li et al. [24] introduced the Bi-Domain Adversarial Neural Network (BiDANN), which accounts for asymmetrical emotional responses in the left and right hemispheres of the brain, leveraging neuroscientific insights to improve emotion recognition performance.
Zhang et al. [34] introduced a cross-subject emotion recognition method that utilizes CNNs with DDC. This method constructs an Electrode-Frequency Distribution Map (EFDM) from EEG signals, using a CNN to extract emotion-related features while employing DDC to minimize distribution differences between source and target domains.
Gokhale et al. [35] proposed an adversarial training approach which learns to generate new samples to maximize exposure of the classifier to the attribute-space, which enables deep neural networks to be robust against a wide range of naturally occurring perturbations.
These deep transfer learning methods has provided valuable insights and strategies for addressing individual differences in EEG signals and achieving significant results in EEG-based emotion recognition. However, a key challenge remains: the dependence of the model on the target domain data during the training process may lead to the model being influenced by the target domain data preference, and increases the practical application cost, which limits the scalability of these models in real-world scenarios.
2.3 Prototype Learning
The core concept of prototype learning is that each class is represented by a prototype (a feature vector that acts as a central, representative feature for that class). Data points belonging to a specific class are clustered around this prototype, enabling classification by evaluating the proximity or similarity of data points to their respective class prototypes. For example,
Snell et al. [36] proposed prototypical networks, which learn a metric space where samples from the same class are clustered around their respective class prototypes.
In Pinheiro et al. ’s work [37], prototype representations are computed for each class, and target domain images are classified by comparing their feature representations to these prototypes, assigning the label of the most similar prototype.
Ji et al. [38] tackled proposed Semantic-guided Attentive Prototypes Network (SAPNet) framework to address the challenges of extreme imbalance and combinatorial explosion in Human-Object Interaction (HOI) tasks.
Liu et al. [39] developed a refined prototypical contrastive learning network for few-shot learning (RPCL-FSL), which combines contrastive learning with few-shot learning in an end-to-end network for enhanced performance in low-data scenarios.
Yang et al. [40] introduced the Two-Stream Prototypical Learning Network (TSPLN), which simultaneously considers the quality of support images and their relevance to query images, thereby optimizing the learning of class prototypes.
These studies show that prototype learning is particularly effective in few-shot learning and unsupervised tasks, significantly improving the robustness of the model and providing a potential approach for the field of affective computing.
Therefore, in the latest research, Zhou et al. [18] proposed a prototype representation based pairwise learning framework (PRPL), which applies prototype learning to EEG emotion recognition, and interacts sample features with prototype features through bilinear transformation.
However, PR-PL considers only class prototypes, assuming that source domain data follow a uniform distribution, a limitation that does not align with real-world variability. In addition, PR-PL still relies on source and target data to align sample features during training, which may limit its applicability in real-world scenarios.
@@ 3 Methodology
The source domain is defined as ={,,,…,, where denotes the number of subjects in the source domain. For each individual subject in the source domain, we have ={,, where denotes the -th sample of -th subject, represents the corresponding emotion label, and is the sample size for the -th subject. The target domain is represented as ={,, where denotes the number of EEG samples in the target domain. For clarity, Tab.1 summarizes the commonly used notations.
@@
Inspired by the work of Peng et al. [41] and Cai et al. [42], we hypothesize that EEG features involves two types of deep features: domain-invariant class features and class-invariant domain features. The domain-invariant class features capture semantic information about the class to which a sample belongs(Such as, negative emotions induce larger LPP amplitudes, and this encoding mechanism is related to individual emotional semantic processing, but does not depend on specific domains.[43]). The class-invariant domain features convey the domain or subject-specific information of the sample(Such as, anatomical differences such as the thickness of the skull and the morphology of the brain regions lead to individual specificity in the spatial distribution of EEG signals. Such characteristics are independent of the emotion class, but affect the baseline of signal amplitude.[44][45]). Overall, original EEG features can be viewed as an integration of these two types of features.
@@
The distributional differences in EEG signals across subjects can be attributed to variations in the domain features, causing a shift in the distribution of class features. Since EEG classification adheres to a common standard, the class features from different subjects should ideally occupy the same feature space and follow a consistent distribution. Traditional methods assume that all test data come from a single, unified domain, effectively focusing only on class features while overlooking the presence of domain features. This assumption limits model generalization across subjects. Feature extraction methods such as DANN can be interpreted as an attempt to remove the domain-specific components from sample features, thereby aligning the class features across subjects and improving the generalization performance of the model.
3.1 Feature Disentanglement
@@
Based on the above theories and assumptions, we consider both domain-specific and class-specific components in EEG feature extraction to improve the robustness and generalization ability across different subjects. To be specific, we start with a shallow feature extractor to obtain shallow features from the EEG samples . Then, we introduce a class feature disentangler and a domain feature disentangler to disentangle the semantic information within these shallow features, resulting in class features and domain features , expressed as:
@@
To improve the effectiveness of the disentanglers in separating the two types of features, we introduce a domain discriminator and a class discriminator . The domain discriminator is designed to determine the domain of the input features, while the class discriminator ascertains the class of the input features. Our goal is for the domain discriminator to accurately identify the domain of the input when given domain features, while the class discriminator should be unable to identify the class based solely on domain features. This inability to classify based on domain features indicates successful disentanglement, where domain features contain only domain-specific information and no class-related information. Similarly, class features only contain class-related information, without domain-specific information.
To achieve this, before the class features are passed into the domain discriminator and the domain features into the class discriminator, they pass through a Gradient Reversal Layer (GRL) to facilitate adversarial training. We use a binary cross-entropy loss function to optimize the discriminators. The output from each discriminator is first passed through a sigmoid layer to obtain probability values, which are then compared to the true labels. This approach converts the multi-class problem into several independent binary classification tasks. The binary cross-entropy loss function is defined as follows:
@@
where represents the GRL. represents the true class labels of -th sample, represents the class feature of -th sample, and represents the domain feature -th sample. This loss function is iteratively optimized during model training to help the class discriminator accurately distinguish the class labels, ensure that the class features only retain class-related information, and minimize the interference of domain-specific features. Similarly, for the domain discriminator, We define the binary cross-entropy loss function as:
@@
here, represents the true domain labels of -th sample. This loss function is optimized to ensure that the domain discriminator accurately identifies the domain-specific information, encouraging the domain features to be disentangled from class-related information. By applying the GRL before the domain features enter the class discriminator, adversarial training is facilitated, further promoting the separation of domain and class features. In the implementation, the shallow feature extractor, domain feature disentangler, and class feature disentangler are all designed as multi-layer perceptrons (MLPs). The disentangled class features and domain features are then utilized in the subsequent prototype inference module. This module learns domain prototypes to represent each domain, and learns class prototypes to capture each class within those domains.
3.2 Prototype Inference